OpenStudy (anonymous):

Debbie Anderson cuts a piece of wood measuring 2 3/8 inches into two pieces. How long is each piece? I got the answer of 19/16 but it needs to be simplified.

OpenStudy (mertsj):

\[\frac{19}{16}=1 \frac{3}{16}\]

OpenStudy (anonymous):

you are the pellet!!

OpenStudy (mertsj):

Whatever that is.

OpenStudy (anonymous):

lol! it wouldnt let me type Sugar Honey Ice Tea

OpenStudy (anonymous):

Yehhh, you learned it now, smitter! 19/8 *2 = 19/16 = 1 3/16

OpenStudy (anonymous):

Can someone explain how the answer was found? Math is not my best subject.